OpenStudy (anonymous):

which one is scalar product? 1.work 2.mass 3.density 4.torque

OpenStudy (anonymous):

Work is scalar product of Displacement and Force . Here is the equation \[W = F \times S \cos \theta \]

OpenStudy (anonymous):

work

OpenStudy (anonymous):

what u did @devarsh.ruparelia??

OpenStudy (anonymous):

dot product is scalar product so none is a scalar product

OpenStudy (anonymous):

scalar product doesn't mean that product of 2 scalars but multiplication of 2 vectors but product gives u scalar work is a vector so it is not scalar product mass as it is not a product of 2 vectors it is not scalar product same with density at last it is torque is a vector so it is not scalar product

OpenStudy (anonymous):

@kr7210 What I did was dot product of 2 vectors which results in the scalar quantity.